OpenStudy (anonymous):

MEDAL ANYONE????? Write the general form of the equation of a sinusoidal function having an amplitude of 2 and a period of Pi/4

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@Callisto

OpenStudy (jango_in_dtown):

period?

OpenStudy (anonymous):

Pi/4

OpenStudy (jango_in_dtown):

phase shift is given?

OpenStudy (anonymous):

no thats all it says

OpenStudy (anonymous):

@TheEdwardsFamily

OpenStudy (anonymous):

@texaschic101

OpenStudy (jango_in_dtown):

The general form is y = Asin(Bx + C) + D whereA is the amplitude,B=2pi/ time period The phase shift of the function is: C/B 'D' is the amount of vertical displacement, or 'y' shift. Here we will keep C and D as it is so A=2 given B=2pi/T=(2pi)/(pi/4)=8 So the genral form is y=2sin(8x+C)+D

OpenStudy (anonymous):

but there is no A B C or D

OpenStudy (jango_in_dtown):

y=2sin(8x+C)+D

OpenStudy (anonymous):

yeah with the answer there is no A,B,C, or D

OpenStudy (jango_in_dtown):

here the phase shift and y shift are not given so you may take it y=2sin 8x

OpenStudy (anonymous):

thank you

OpenStudy (jango_in_dtown):

A, B,C,D we calculate from the given data.... A is the amplitude, B is 2pi/timeperiod.

OpenStudy (anonymous):

oh